### v1

# XATU: A Fine-grained Instruction-based Benchmark for Explainable Text Updates

Abstract
Text editing is a crucial task that involves modifying text to better align with user intents. However, existing text editing benchmark datasets have limitations in providing only coarse-grained instructions. Consequently, although the edited output may seem reasonable, it often deviates from the intended changes outlined in the gold reference, resulting in low evaluation scores. To comprehensively investigate the text editing capabilities of large language models, this paper introduces XATU, the first benchmark specifically designed for fine-grained instruction-based explainable text editing. XATU covers a wide range of topics and text types, incorporating lexical, syntactic, semantic, and knowledge-intensive edits. To enhance interpretability, we leverage high-quality data sources and human annotation, resulting in a benchmark that includes fine-grained instructions and gold-standard edit explanations. By evaluating existing open and closed large language models against our benchmark, we demonstrate the effectiveness of instruction tuning and the impact of underlying architecture across various editing tasks. Furthermore, extensive experimentation reveals the significant role of explanations in fine-tuning language models for text editing tasks. The benchmark will be open-sourced to support reproduction and facilitate future research. [^1] The code will be available at https://github.com/megagonlabs/xatu
1 Introduction
Figure 1: Illustrated examples of coarse- and fine-grained instructions for text editing. LLMs can accurately perform text editing based on coarse-grained instructions, but may not meet the user’s intention. In contrast, fine-grained instructions lead to accurate and user-intended text editing.
Recent advances in large language models (LLMs) have demonstrated remarkable zero-shot text generation capabilities across a wide range of downstream natural language processing (NLP) tasks like question answering, dialogue, and summarization Radford et al. (2019); Raffel et al. (2020); Brown et al. (2020); Zhang et al. (2022, 2023a). By incorporating instruction tuning Wei et al. (2022) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Ouyang et al. (2022), state-of-the-art LLMs exhibit a remarkable ability to effectively align with user instructions.
However, most such work adopts a one-shot paradigm, generating text in a single step. In contrast, human text composition often involves an iterative process consisting of multiple drafts and revisions Faltings et al. (2021). This iterative process allows for refining and improving the text over time, ensuring better alignment with the user’s intents, while the current setting of generating text contents in a single pass can be highly restrictive Hayes and Flower (1980); Collins and Gentner (1980). An alternative approach that offers more flexibility is iterative text generation and refinement, involving successive iterations of modular updates and modifications to the text Vaughan and McDonald (1986). These iterations focus on text editing tasks such as paraphrasing and updating information to enhance the overall quality of the generated output.
Benmark Domain Reference Instruction Explanation
WikiSemanticIntention Yang et al. (2017) wiki
WikiAtomicEdits Faruqui et al. (2018) wiki
WikiDocEdits Faltings et al. (2021) wiki coarse-grained
ITERATER Du et al. (2022) multiple
EditEval Dwivedi-Yu et al. (2022) multiple coarse-grained
XATU (ours) multiple fine-grained
Table 1: A detailed comparison of XATU with existing text editing datasets and benchmarks.
Consequently, there has been a growing interest among researchers in instruction-based text editing tasks in recent years. Text editing plays a crucial role in enabling interactivity, as it allows users to modify previously generated text to suit their dynamic constraints. This represents a step further beyond one-shot generation towards interactive text generation. Specifically, Faltings et al. (2021) introduced the concept of controllable text editing, wherein a model learns to edit the text in response to user commands while incorporating grounding to prevent hallucination issues. Moreover, Dwivedi-Yu et al. (2022) proposed EditEval, a benchmark and evaluation suite that utilizes high-quality existing and new datasets to automatically assess the editing capabilities of different systems.
Nevertheless, these works only consider simple commands or instructions, such as ‘Update the article’ or ‘Expand’, resulting in three main limitations. (a) These instructions are often too coarse-grained, to the extent that even humans may struggle to understand the desired intent of the edit, limiting models’ instruction following capabilities. (b) It poses challenges for text editing evaluation since outputs obtained using coarse-grained instructions may not accurately reflect the true editing capabilities of LLM systems. The coarse-grained instructions cannot provide sufficient guidance for the editing system, so the edited output may appear correct but deviate from the intended changes, as shown in Figure 1. The edited output updates the death information of Isidore Mankofsky correctly but since the ground truth output updates the information about his famous works. This edit will receive very low evaluation scores since current automatic evaluation methods for text editing systems are mostly based on surface-level overlap with human-written gold references. (c) The current text editing systems suffer from a lack of explanation ability, which in turn limits the interpretability of the edits made. This lack of interpretability hampers our understanding of the underlying text editing behaviors.
To address the aforementioned limitations, we present XATU, a novel fine-grained instruction-based benchmark designed for explainable text updates. The benchmark leverages high-quality existing data sources to enable automatic evaluation of LLM editing capabilities. In comparison to other datasets and benchmarks, XATU highlights its inclusion of a wider range of diverse topics, fine-grained edit instructions, and corresponding explanation rationales, as shown in Table 1. By utilizing XATU, we are able to evaluate and compare the performance of several state-of-the-art language models, considering both zero-shot and fine-tuning settings. Experiment results emphasize the significant role of explanations in the fine-tuning process of language models for text editing tasks. Our findings shed light on the importance of incorporating explanation rationales to enhance the performance of language models in text editing, ultimately leading to improved output quality. We summarize the contributions of this paper as follows:
  • We propose XATU, a new benchmark for instruction-based text editing. We collect high-quality data from a set of tasks and datasets for iterative text updates and provide gold fine-grained instructions and explanations with both LLM generation and human annotations. To the best of our knowledge, XATU is the first text editing dataset with fine-grained instructions and explanations.
  • We conduct a thorough evaluation of existing open and closed large language models on our benchmark with different input formats to test their text editing capabilities. We further investigate the influence of instruction fine-tuning using explanations and fine-grained editing instructions.
2 Related Work
2.1 Text Editing
Several prior studies have concentrated on the domain of text editing. Guu et al. (2018) introduced a method that involves generating sentences through the editing of training prototypes. Faruqui et al. (2018) presented the WikiAtomicEdits dataset, which comprises edits extracted from Wikipedia editing history. Subsequently, Yin et al. (2019); Reid and Neubig (2022) leveraged Wikipedia editing history to propose the task of learning to represent edits. Furthermore, Iso et al. (2020) introduced a fact-based text editing task in their work.
Recently, text editing has been approached as an interactive task, leading to the development of command-based editing systems Faltings et al. (2021) and interactive editing systems Schick et al. (2023). Researchers have also explored the integration of text editing into interactive self-refinement frameworks for text generation using large language models. Welleck et al. (2023) introduced a self-corrective learning framework that incorporates a corrector into the language model, enabling self-correction during sequence generation. Akyürek et al. (2023) proposed a reinforcement learning-based approach for generating natural language feedback to correct generation errors. Furthermore, Zhang et al. (2023b) presented an iterative refinement framework for abstractive summarization.
2.2 Text Editing Datasets
Previous research has also introduced various datasets and resources that focus on iterative text revisions within specific domains. For instance, datasets presented in  Yang et al. (2017); Faruqui et al. (2018); Anthonio et al. (2020) primarily concentrated on the domain of Wikipedia edit history, while Spangher et al. (2022) developed a dataset tailored to news articles. Researchers in the field have also proposed more comprehensive benchmarks that cover multiple domains. IteraTeR Du et al. (2022) provides iterative tasks from multiple domains, but has only a limited number of tasks, such as fluency, coherence, clarity, style, and meaning change. EditEval Dwivedi-Yu et al. (2022) is perhaps the closest to our benchmark XATU in that it covers data from multiple domains: Wikipedia, Wikinews, news articles, and arXiv.
Moreover, it is important to note that only EditEval and WikiDocEdits Faltings et al. (2021) provide short and simple instructions or commands to guide the editing process. In contrast, our benchmark XATU distinguishes itself by including fine-grained instructions for each data instance, which is necessary for assessing the genuine text editing capabilities of different systems. Furthermore, XATU also includes the rationale explanations for each of its instances, which have been proven to be significant for instruction fine-tuning models Hsieh et al. (2023); Mukherjee et al. (2023).
3 The XATU Benchmark
3.1 Data Source
In line with the systematic formulation of natural language styles in natural languages presented by DiMarco and Hirst (1993), we adopt a similar approach to categorize text editing into four main editing aspects: lexical, syntactic, semantic, and knowledge. Our benchmark XATU aims to serve as a relatively comprehensive benchmark and a holistic evaluation framework for text editing tasks, covering all four aspects of text editing.
To ensure the comprehensiveness and quality of our benchmark, we meticulously curate high-quality editing data from $4$ distinct downstream NLP tasks, encompassing a total of $1000$ annotated data from $9$ data sources, inspired by the criteria outlined in DiMarco and Hirst (1993). A summary of the tasks and datasets, along with an overview of the editing aspects covered by each dataset, is presented in Table 2. We also transform all datasets within the XATU benchmark into a consistent format, extracting the input text, gold edits, task type, and optional reference documents from the following data sources, as illustrated in the example in Figure 2. This standardized format facilitates evaluation and comparison across different datasets within the benchmark.
{tcolorbox}
[fontupper=, colback=green!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Instruction: Modify the input text to incorporate details about the nature of the matches played, the players included in the tour, mention of Nat Sciver captaining the England team for the first time, and the simultaneous scheduling of New Zealand men’s matches. {tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Input: The England women’s cricket team is scheduled to play against New Zealand women’s cricket team in February and March 2021. The tour is scheduled to be played during the time that was originally scheduled to be used to host the … {tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Reference: (0) The last three T201 matches took place on the same day as the New Zealand women’s fixtures at the same venues…
(7) Sophie_Ecclestone International career In February 2021, during England’s tour of New Zealand, Ecclestone took her 100th international wicket.
(8) Sarah_Glenn International career In 2021, Glenn was named in the squad for England’s tour of New Zealand.
{tcolorbox}[fontupper=, colback=blue!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Output: The England women’s cricket team played against the New Zealand women’s cricket team in February and March 2021. The series was played during the time that was originally scheduled to be used to host the 2021 Women’s Cricket World Cup, after the tournament was postponed by one year due to the COVID-19 pandemic… {tcolorbox}[fontupper=, colback=green!10, left=1mm, right=1mm, top=1mm, bottom=1mm] Explanation: The text editing system was instructed to update the text based on the references provided. It corrected the phrasing to past tense, indicating that the England and New Zealand women’s cricket teams already ’played’ the matches instead of ’scheduled to play’. The information about the tour, including the confirmation of fixtures…
Figure 2: The instance format of the data in XATU benchmark. Data in blue (Input, Output, Reference) are extracted from the corresponding data sources, and data in green (Fine-grained instruction and explanation) are obtained from joint automatic and human annotations.
Grammar Error Correction
The first task in XATU is to identify and correct errors that don’t follow grammar rules. We first incorporate data from JHU FLuency-Extended GUG (JFLEG) Napoles et al. (2017), a dataset specifically designed for the task of grammar error correction. JFLEG uses holistic fluency edits to not only correct grammatical errors but also make the original text more native sounding. By including this dataset, we aim to assess the basic ability of text editing systems to improve fluency and grammatical accuracy simultaneously.
Simplification
We also include ASSET Alva-Manchego et al. (2020), a text simplification dataset that provides manually produced simplifications through a wide range of transformation techniques. We aim to evaluate the text editing capabilities of systems when it comes to simplifying complex texts while preserving their essential meaning.
Style Transfer
Text editing also involves transferring the styles of the input sentences, so we include the following three datasets, aiming to evaluate the effectiveness of text editing systems in handling style transfer and bias mitigation tasks.
Task Dataset Train Test Aspect
Grammar JFLEG 10 50 Lexical, Syntactic
Simplification ASSET 10 50 Lexical, Syntactic
Style Transfer WNC 20 100 Lexical, Semantic
Wikibias 20 100
StylePTB 10 50
Information Update FRUIT 30 150 Semantic, Knowledge
Evidence 40 200
DeFacto 30 150
Factedit 30 150
Table 2: The detailed datasets included in the XATU benchmark.
We first include the Wiki Neutrality Corpus (WNC) Pryzant et al. (2020), a collection of original and de-biased sentence pairs mined from Wikipedia edits by carefully filtering based on the editor’s comments.
We also adopt data from Wikibias Zhong (2021), a manually annotated parallel corpus with sentence pairs from Wikipedia editing history. The inclusion of Wikibias complements the mined data from WNC, providing additional annotations to address both sentence-level and token-level biases.
Furthermore, we also include data from StylePTB Lyu et al. (2021). StylePTB contains paired sentences undergoing various fine-grained stylistic changes and compositions of multiple transfers.
Information Update
Our XATU benchmark also includes the following four knowledge-intensive text editing datasets. These datasets are particularly challenging as they require text editing systems to update the input text based on given instructions and external reference evidence. These knowledge-intensive datasets offer a valuable assessment of the systems’ capacity to leverage external information for text editing tasks.
We include the FRUIT dataset (Iv et al., 2022), a dataset collected by comparing two snapshots of the same Wikipedia article. The reference documents were identified by searching for other Wikipedia articles and human filtering. We include this gold set from the FRUIT dataset in XATU.
Evidence Thorne and Vlachos (2021) is a dataset created using a two-stage distant supervision approach, where evidence is incorporated into masked claims from the FEVER dataset Thorne et al. (2018). Each claim in the dataset is paired with reference evidence claims obtained from Wikipedia.
DeFacto Liu et al. (2022) dataset contains consists of document summaries from the XSum dataset Narayan et al. (2018), accompanied by human-corrected versions that rectify factual errors present in the original summaries. For the text editing task, we utilize the original document as the reference to evaluate the ability of text editing systems to correct factual errors in the summaries effectively. The inclusion of this dataset allows us to assess the performance of systems in editing text to ensure factual accuracy, an important aspect of text editing in real-world applications.
FactEdit Iso et al. (2020) is an editing dataset that incorporates facts sourced from a knowledge base, such as multiple triples, as the reference. This dataset is constructed from table-to-text data sources.
Previous work has explored the use of Hamming distance for sentence-level editing task difficulty estimation Lyu et al. (2021). However, as our input and output texts may have different lengths, we employ a token-level Levenshtein distance to measure the relative difficulty of the different datasets in our benchmark. Based on this distance measurement, we categorize the tasks into three levels: easy, medium, and hard. The results of this categorization can be found in Table 3. We notice that information update tasks generally require more editing operations compared to simple lexicon tasks like grammar error correction and neutralization. This highlights the increased complexity and difficulty associated with tasks that involve updating and incorporating external information into the text. To ensure that the benchmark presents a challenging evaluation environment, we deliberately incorporate more examples from demanding tasks.
[…]
Easy JFLEG 1.47
WNC 1.58
STYLEPTB 1.72
Wikibias 1.92
Medium Evidence 3.56
ASSET 4.72
DeFacto 4.82
Hard Factedit 9.75
FRUIT 13.62
Table 3: Average token-level Levenshtein distance between the input and edited output in XATU.
3.2 Annotation Process
Serving as a fine-grained instruction-based and explainable benchmark, each instance in XATU comprises five key components as shown in Figure 2: inputs, outputs, optional references, fine-grained instructions, and explanations. Since the source datasets lacked fine-grained instructions and explanations, we meticulously annotated them.
Challenges in Crowdsourcing
Annotating fine-grained instructions and explanations poses a challenge due to the inherently open-ended nature of the text production task, making it difficult to filter out low-quality workers. Additionally, many workers turn to LLMs for text production tasks due to their prevalence Veselovsky et al. (2023), but the output quality heavily depends on the user’s LLM skills Zamfirescu-Pereira et al. (2023). Therefore, relying solely on crowd workers for annotation may not be advisable.
To address these concerns and maintain the quality and reliability of the annotated data, we implemented an LLM-in-the-loop annotation approach, which involves generating candidates of fine-grained instruction and explanations using LLMs and then validating the output with human workers. This allowed us to mitigate the challenges associated with crowd-sourcing and refine our methodology for generating and evaluating fine-grained instructions and edit explanations.
{tcolorbox}
[fontupper=, fontlower=, colback=gray!5, left=1mm, right=1mm, top=1mm, bottom=1mm] Input: At least once an episode we see protestors marching around screaming slogans
[…]
Output: <ins>One or more times</ins> an episode protestors <ins>are seen walking</ins> around <ins>yelling blue</ins> slogans <ins>.</ins>
Figure 3: Illustrated example of adding HTML tags to the input and output to explicitly indicate the edited portion to LLMs.
Candidate Generation by LLMs
We generate candidates of fine-grained instructions and explanations for each instance using LLMs Zhou et al. (2023b); Honovich et al. (2023). For the LLMs to generate candidates, we use GPT-4 OpenAI (2023). Through our preliminary prompt engineering, we found that merely asking LLMs to generate the instructions and explanations often resulted in undesired outputs [^2] More details can be found in the Appendix.. We found that adding HTML tags to explicitly indicate the edited portion is highly effective, as illustrated in Figure 3. Thus, we decided to use this prompting technique to generate instructions and explanations. The full prompt can be found in the Appendix.
Note that our analysis showed no significant differences between coarse and fine-grained instructions for easy tasks (i.e., JFLEG, WNC, STYLEPTB, and WikiBias) since the inputs are relatively short and the edit distance is relatively small. Consequently, we chose to only conduct fine-grained instruction annotation on the more complex information update task ($65\%$ data).
Candidate Validation by Human
We conducted candidate validation for fine-grained instructions and explanations by employing human evaluation through the Appen platform. [^3] https://appen.com/ Workers were instructed to assess the validity of generated instructions or explanations based on input, output, and optional references. To ensure quality, we selected and manually annotated $20$ random test examples and accepted workers who provided correct answers for over $80\%$ of them. If a candidate was considered invalid, we repeated the candidate generation process for that instance.
We evaluated the quality of annotations through inter-annotator agreement, which resulted in an overall agreement rate of 80.27%. We annotated the dataset based on difficulty levels defined in Table 3. A detailed breakdown of the number of instances annotated is provided in Table 2. The result of this rigorous process was the creation of the XATU benchmark, containing 1000 high-quality, fine-grained text editing instances.
3.3 Benchmark Usage
With the inclusion of fine-grained instructions and corresponding editing rationale explanations, our benchmark offers comprehensive support for a range of traditional text editing tasks like edit representation modeling, automatic editing instruction generation, and editing span prediction. Additionally, the unique structure of XATU enables two new tasks: editing explanation generation and editing evidence retrieval. Our annotated edit explanations can be used to train and assess models aimed at generating coherent and informative explanations for the editing process. Moreover, the information update data that incorporates external reference documents allows for the development and evaluation of models that can effectively retrieve the necessary evidence for text editing tasks.
By providing support for these downstream tasks, our benchmark aims to facilitate advancements in various aspects of text editing, fostering the development of models and techniques that can improve the efficiency and effectiveness of text editing systems.
4 Experiments
In this section, we perform a series of text editing experiments using the XATU benchmark. These experiments aim to investigate and address the following research questions:
  • Q1: What is the actual text editing ability of existing open/closed large language models?
By evaluating the performance of LLMs, we aim to assess their effectiveness in carrying out text editing tasks and understand their limitations and strengths in this context.
  • Q2: How do our fine-grained instructions differ from the coarse-grained instructions that are commonly provided in text editing tasks?
By comparing the performance of models using fine-grained instructions from XATU, we seek to highlight the impact of instruction granularity on the quality and accuracy of text editing.
  • Q3: How do the explanations provided in XATU benefit model tuning under the fine-tuning settings?
By examining the performance of models that incorporate the explanations provided in XATU during fine-tuning, we aim to evaluate the effectiveness of these explanations in improving model performance and understanding the role of explanations in the context of text editing.
Model Setting JFLEG ASSET WNC Wikibias PTB FRUIT Evidence DeFacto Factedit
Flan-T5 coarse 64.98 52.01 62.54 55.58 51.05 19.06 39.13 36.77 7.98
fine - - - - - 16.84 41.84 42.87 6.80
Exp. 59.09 46.47 56.71 51.98 46.78 21.40 53.34 38.78 11.65
Flan-UL2 coarse 64.35 50.08 59.39 51.88 51.83 19.84 52.61 31.15 23.06
fine - - - - - 43.38 68.79 52.54 35.15
Exp. 83.38 66.82 86.78 79.22 74.81 38.08 74.35 52.83 29.13
Alpaca coarse 68.14 41.20 45.68 42.35 54.14 51.75 53.65 52.58 41.08
fine - - - - - 52.74 69.42 60.81 34.83
Exp. 75.82 62.41 70.90 69.53 77.99 54.00 74.12 66.17 39.89
GPT3 coarse 50.74 30.82 32.24 34.58 42.63 26.47 34.12 28.42 31.46
fine - - - - - 27.43 36.72 34.85 36.52
Exp. 56.34 37.42 41.34 42.48 47.23 32.54 47.23 37.28 38.75
GPT4 coarse 70.32 56.71 64.18 58.13 58.72 43.28 58.19 54.62 43.59
fine - - - - - 49.28 62.34 62.42 48.27
Exp. 84.58 73.54 89.23 82.93 84.39 59.43 81.38 71.38 58.38
Table 4: Zero-shot text editing results of different LLMs under different prompt settings on XATU benchmark. We omit results on simple datasets that do not require fine-grained instruction. We use coarse, fine, and Exp. to represent text editing with coarse-grained instructions fine-grained instructions, and explanations, respectively.
4.1 Experimental Setup
We include the following baseline large languages models in our experiments:
  • GPT-3 (Brown et al., 2020) is a 175B parameter pre-trained decoder-only model.
  • GPT-4 (OpenAI, 2023) is the most recent multimodal large language model created by OpenAI. It was pre-trained to predict the next token and was then fine-tuned with reinforcement learning from human and AI feedback. We evaluate both GPT-3 and GPT-4 through OpenAI’s API. [^4] https://beta.openai.com/
  • T5 Raffel et al. (2020) is an encoder-decoder model pre-trained on a multi-task mixture of unsupervised and supervised tasks.
  • Flan-T5 Chung et al. (2022) is an enhanced version of T5 that has been instruction fine-tuned on a mixture of tasks.
  • UL2 Tay et al. (2023) is a unified framework for pre-training models with Mixture-of-Denoisers, a pre-training objective that combines diverse pre-training paradigms together.
  • Flan-UL2 is an encoder-decoder model that uses the same configuration as the UL2 model. It was fine-tuned using the "Flan" prompt tuning and dataset collection.
  • LLaMa Touvron et al. (2023) is a collection of foundation language models trained on trillions of tokens. It is an auto-regressive language model built on transformer architecture.
  • Alpaca Taori et al. (2023) is an instruction-following language model fine-tuned from the LLaMA 7B model on 52K instruction-following demonstrations.
Note that PEER Schick et al. (2023) is another strong text editing baseline according to results in Dwivedi-Yu et al. (2022). We didn’t include it since the model checkpoint is not publicly available.
For evaluation, we use SARI scores Xu et al. (2016), an $n$-gram based metric commonly used for measuring editing tasks such as simplification Zhao et al. (2018) and sentence fusion Malmi et al. (2019). It has been demonstrated to correlate most closely with human judgment and is computed as:
$SARI=(F1_{add}+F1_{keep}+P_{del})/3,$
where $F1_{add},F1_{keep},P_{del}$ represent the F1 scores and precision for add, keep, and delete operations, respectively. We utilize the Hugginface implementation of SARI  [^5] https://huggingface.co/spaces/evaluate-metric/sari, where $n=4$. EditEval Dwivedi-Yu et al. (2022) includes more n-gram-based valuation metrics, but the results do not demonstrate any significant difference from SARI scores.
Model Setting JFLEG ASSET WNC Wikibias PTB FRUIT Evidence DeFacto Factedit
T5 coarse 62.75 52.26 64.20 56.01 56.59 45.27 60.03 59.07 47.65
fine 63.43 51.85 58.65 56.93 61.07 48.42 71.56 60.91 45.80
Exp. 72.73 60.23 77.39 70.76 72.22 43.83 77.24 64.23 46.14
Flan-T5 coarse 64.07 52.71 65.04 58.86 63.63 50.95 62.81 59.40 45.99
fine 65.39 53.00 65.11 59.56 63.15 53.64 76.43 66.62 47.52
Exp. 79.17 69.18 84.67 75.31 75.64 52.33 85.60 71.25 47.90
LLaMA coarse 63.86 45.46 62.84 53.72 58.87 49.18 63.69 52.09 50.25
fine 66.18 47.56 64.30 61.08 60.67 53.44 82.19 69.38 52.45
Exp. 83.31 70.28 91.22 84.66 84.54 54.22 86.85 79.14 55.45
Alpaca coarse 65.71 44.95 63.68 55.77 63.62 49.18 64.13 56.73 46.73
fine 69.64 47.14 62.78 50.57 61.18 51.81 83.48 73.69 46.35
Exp. 83.52 70.83 87.93 75.85 83.56 58.33 88.41 77.91 47.45
UL2 coarse 66.61 54.55 69.82 60.45 61.86 51.49 70.65 59.21 58.18
fine 71.22 54.84 71.19 56.04 67.27 56.58 84.37 70.27 56.06
Exp. 87.81 78.22 91.79 82.24 88.10 54.64 90.54 78.65 56.44
Flan-UL2 coarse 68.03 52.34 73.93 57.24 73.12 51.81 71.80 58.01 58.72
fine 71.84 52.59 75.16 58.09 69.38 63.91 86.17 73.56 60.64
Exp. 90.44 79.84 94.68 86.23 86.06 59.46 91.71 82.84 60.76
Table 5: Instruction fine-tuning results of different LLMs under different prompt settings on XATU. We use coarse, fine, and Exp. to represent text editing with coarse-grained instructions fine-grained instructions, and explanations, respectively.
Implementation Details
We use the Huggingface implementation of all models with LoRA Hu et al. (2022) for the fine-tuning experiments. We instruction-tuned all models for $15$ epochs with $200$ training examples at a learning rate of $51e-$ 5 and tested on the original $1000$ test set, as shown in Table 2. We keep a small training data size as Zhou et al. (2023a) found that alignment doesn’t require too much data. The detailed training parameters are presented in Section B.
4.2 Zero-shot Results
The zero-shot evaluation results on the XATU benchmark are presented in Table 4, which highlights several important observations.
GPT-4 demonstrates exceptional zero-shot editing performance, surpassing all other large language models by a significant margin across all datasets. Alpaca emerges as the second-strongest model, but there is still a considerable performance gap when compared to GPT-4. On the other hand, models without instruction tuning, such as GPT-3, struggle to follow the instructions effectively, resulting in relatively low SARI scores. This emphasizes the importance of instruction tuning in enhancing the text editing performance of language models.
In addition, we find that almost all models exhibit improvements when guided by the fine-grained instructions provided in XATU, as opposed to simple and coarse-grained instructions. This underscores the significance of fine-grained instructions in enhancing the performance and accuracy of text editing tasks. Notably, the inclusion of explanations as guidance during the text editing process leads to further improvements across all models.
Furthermore, our findings reveal that the underlying architecture (encoder-decoder vs. decoder-only) of language models significantly impacts the performance of different types of text editing tasks. We can see that Flan-T5 and Flan-UL2 exhibit biased performance, excelling in more straightforward tasks like style transfer but facing challenges in more complex tasks like information updates. In contrast, the Alpaca model performs well on information update tasks but achieves lower scores in simple neutralization tasks. We attribute this difference in performance to the underlying architecture of the foundation models: Flan-T5 and Flan-UL2 are both encoder-decoder models, while Alpaca is built upon the LLaMa decoder-only model. The results demonstrate that the encoder’s ability to understand and represent the input text is more important for lexical and syntactic editing, while the decoder’s capacity to generate new and relevant text appears to be more influential for knowledge-intensive tasks.
4.3 Fine-tuning Results
After fine-tuning all models using different prompting formats, we evaluate the fine-tuned models on XATU, and the results are presented in Table 5. A notable observation is the significant performance improvement compared to the zero-shot results across all datasets and under all settings. This demonstrates the effectiveness of instruction fine-tuning, even with a limited number of examples for text editing tasks ($200$ in this case).
Among all the instruction fine-tuned models, Flan-UL2 consistently exhibits the strongest performance across all three settings. This indicates its strong adaptation capability during the fine-tuning process. Furthermore, we observe that the fine-grained instructions in XATU lead to larger performance improvements in tasks that have fine-grained instructions (e.g. FRUIT, DeFacto), while the improvements are comparatively smaller in simpler tasks that lack fine-grained instructions (e.g. WNC, PTB) in the fine-tuning data.
Comparing T5 and UL2 with their instruction-tuned versions (Flan-T5 and Flan-UL2), we observe that the instruction-tuned versions of LLM consistently outperform their base models. This is especially evident when the models are guided with fine-grained instructions or explanations. These results further emphasize the effectiveness of the instruction alignment ability that these models acquire through instruction fine-tuning.
Overall, the fine-tuning process significantly improves the performance of the models on text editing tasks. The presence of fine-grained instructions and explanations further enhances the performance.
4.4 Discussion
Figure 4: Boxplot comparing improvements of instruction fine-tuning vs. zero-shot performance.
Figure 5: Boxplot comparing instruction-tuned model vs. pre-trained counterparts.
As illustrated in the left side of Figure 4, we compare the results of three models after instruction fine-tuning with their zero-shot performance. We observe that the fine-tuned models consistently outperform their zero-shot counterparts, indicating the effectiveness of instruction fine-tuning in improving text editing performance. Across both coarse-grained and fine-grained instruction settings, we find that the Flan-UL2 model achieves the largest average improvements. On the other hand, Alpaca demonstrates superior robustness across all text editing tasks compared to Flan-T5 and Flan-UL2. This enhanced robustness may be attributed to the underlying decoder-only architecture.
As depicted in the right side of Figure 4, the Flan-UL2 model consistently outperforms the UL2 model across all settings. Similarly, the Flan-T5 model performs better than the base T5 model. These results highlight the effectiveness of instruction tuning in improving the performance of language models across different settings. The instruction-tuning procedure allows the models to better understand and follow the provided instructions, resulting in more accurate and appropriate text edits. We also notice edits generated with coarse-grained instructions show better robustness across all tasks.
5 Conclusion
This paper introduces XATU, the first benchmark for explainable text updates with fine-grained instructions. XATU is a diverse benchmark covering a wide range of topics and text types and leverages high-quality data sources from various existing sources. We compare existing open and closed instruction-tuned language models under both the zero-shot and fine-tuning settings and real their capabilities to edit text and follow instructions. By releasing the benchmark to the community, we hope to stimulate further research in developing instruction-based text editing models and even potentially interactive text generation systems.
Limitations
Our benchmark XATU still has several limitations. First, XATU only includes data sources from the English language. We leave evaluating multilingual text editing capabilities to future work.
Second, XATU only has thousands of examples due to budget limitations. However, experiment results demonstrate that instruction fine-tuning with $200$ high-quality examples already shows significant improvements.
Third, we only compare LLMs performance under zero-shot and fine-tuning settings. We leave in-context learning results (few-shot) to future work.
Appendix A Data Annotation
The human annotations were conducted on the Appen [^6] https://appen.com/ platform, an example user interface is shown in Figure 6.
Figure 6: The interface used for annotation.
The prompts used to generate the explanation are:
{tcolorbox}
[fontupper=, title=Prompts to generate fine-grained instruction] [no prose]
# Task:
Your task is to write a detailed instruction that enables the AI assistant to edit the original text into a revised text based on the references. The instruction must not cause information leakage about the revised text.
# Original text:
{{input}}
# Reference:
{{reference}}
# Revised text:
{{output}}
# Instruction:
{tcolorbox}
[fontupper=, title=Prompts to generate explanation] [no prose]
# Task:
Your task is to provide a two-sentence explanation of the edits made based on the instruction and reference by comparing the original and revised texts.
# Instruction:
{{instruction}}
# Original text:
{{input}}
# Reference:
{{reference}}
# Revised text:
{{output}}
# Explanation:
Appendix B Implementation Details
The detailed instruction fine-tuning parameters are summarized in Table 6. We use LoRA Hu et al. (2022) to instruction fine-tune all models for $15$ epochs with $200$ additional training examples at a learning rate of $51e-$ 5.
$\mathtt{lora}_{\mathtt{r}}$ 8
$\mathtt{lora}_{\alpha}$ 32
$\mathtt{num\_epochs}$ 15
lr 1e-5
$\mathtt{lora}_{\mathtt{dropout}}$ 0.1
$\mathtt{seed}$ 634
$\mathtt{max\_length}$ 1024
Table 6: Hyper-parameters used in the fine-tuning experiments.
The prompts used to generate the text editing and fine-tuning are:
{tcolorbox}
[fontupper=, title=Prompts for Text Editing] [no prose]
Below is an instruction that describes a task, along with an input text paired with a reference and an explanation that provide further context. Please edit the input text based on the instruction, the reference, and the explanation. Your response should only include the edited output.
# Instruction:
{{instruction}}
# Input:
{{input}}
# Reference:
{{reference}}
# Explanation:
{{explanation}}
# Response: